### v2

# Harmonic maps from $\mathbb{S}^{3}$ to $\mathbb{S}^{2}$ and the rigidity of the Hopf fibration

Abstract.
It was conjectured by Eells that the only harmonic maps $f:\mathbb{S}^{3}\to\mathbb{S}^{2}$ are Hopf fibrations composed with conformal maps of $\mathbb{S}^{2}$. We support this conjecture by proving its validity under suitable conditions on the Hessian and the singular values of $f$. Among the results, we obtain a pinching theorem in the spirit of that of Simons, Lawson and Chern, do Carmo and Kobayashi for minimal hypersurfaces in the sphere.
Key words and phrases:
Harmonic maps, weakly conformal maps, Hopf fibration
2000 Mathematics Subject Classification:
Primary 53C43, 58E20, 53C24, 53C40, 53C42, 57K35
A. Georgakopoulos & A. Savas-Halilaj are supported in the framework of H.F.R.I. call “Basic Research Financing” under the National Recovery and Resilience Plan “Greece 2.0” funded by the European Union-NextGenerationEU (HFRI Project Number: 14758). L. Mari is supported by the PRIN 20225J97H5 “Differential-geometric aspects of manifolds via Global Analysis”
1. Introduction
A smooth map $f:M\to N$ between compact Riemannian manifolds is called harmonic if it is a critical point of the energy functional
$\mathcal{E}(f)\doteq\frac{1}{2}\int_{M}|df|^{2}\,d\mu,$
where $|df|$ denotes the norm of the differential $df$ with respect to the Riemannian metrics on $M$ and $N$, and $d\mu$ is the volume form of $M$. A classical question in the theory of harmonic maps is whether every continuous map admits a harmonic representative in its homotopy class. Eells and Sampson [eells] introduced the heat flow to deform a given map into a harmonic one, which leads to a harmonic representative if the target manifold is non-positively curved, but when $N$ has positive curvature the existence problem is more subtle. In particular, the classification of harmonic maps between spheres is far from being completely understood.
In 1931 Hopf [hopf] constructed a remarkable example of a homotopically nontrivial harmonic map $\pi:\mathbb{S}^{3}\to\mathbb{S}^{2}$ now known as the Hopf fibration, see Section 3.1. According to results of Eells, Ferreira and Ratto [ratto2, eells1, eells2], for every homotopy class in $\mathbb{S}^{3},\mathbb{S}^{2}]$ there exist Riemannian metrics on $\mathbb{S}^{3}$ and $\mathbb{S}^{2}$ for which this class admits a harmonic representative. In contrast with the previous results, for the standard round metrics on $\mathbb{S}^{3}$ and $\mathbb{S}^{2}$ the only known examples of harmonic maps $f:\mathbb{S}^{3}\to\mathbb{S}^{2}$ are compositions of the Hopf fibration with conformal maps of $\mathbb{S}^{2}$. Indeed, Eells formulated the following conjecture [baird1, Note 10.4.1, p. 422] and [wangg, p. 730]:
Conjecture (Eells).
Every harmonic map $f:\mathbb{S}^{3}\to\mathbb{S}^{2}$ (with their round metrics) factors through the Hopf fibration as $f=g\circ\pi,$ where $g:\mathbb{S}^{2}\to\mathbb{S}^{2}$ is a conformal map.
Currently, the conjecture is still widely open. The purpose of the present paper is to support its validity by proving it under some further mild assumptions on $f$. We shall investigate harmonic maps $f:\mathbb{S}^{3}\to\mathbb{S}^{2}$ satisfying suitable bounds on the Hessian $B=\nabla df$ and the singular values $\lambda_{1}\geq\lambda_{2}\geq\lambda_{3}$ of $df$, i.e. the eigenvalues of $\sqrt{df^{*}df}$. Note that, for dimensional reasons, we have that $\lambda_{3}=0$. In particular, the $2$-dilation of $f$ is defined by
$\operatorname{D}_{2}\doteq\lambda_{1}\lambda_{2},$
The quantity $\operatorname{D}_{2}$ measures the area distortion of $f$ on tangent $2$-planes. The study of the $2$-dilation of maps between Riemannian manifolds has recently led to several interesting rigidity and topological applications; see, for example, [savas5, savas4, tsai, lee1, lee2, ding1, ding2].
The first main result of our paper is the following one. In a certain sense, the theorem can be considered as an analogue of the well-known pinching property for the second fundamental form of minimal hypersurfaces in spheres due to Simons [simons], Lawson [lawson] and Chern, Do Carmo and Kobayashi [CdCK].
Theorem A.
Suppose that $f:\mathbb{S}^{3}\to\mathbb{S}^{2}$ is a harmonic map. Then the following hold:
  1. (1)
    If the squared norm of its Hessian $B$ satisfies
    $|B|^{2}\leq\operatorname{D}_{2}(\operatorname{D}_{2}+a),\quad\text{for some}\,\,\,a\in[0,2),$ (1.1)
    then $f$ is the composition of a Hopf fibration with a holomorphic map $g:\mathbb{S}^{2}\to\mathbb{S}^{2}$ (up to conjugation).
  2. (2)
    If the squared norm of its Hessian $B$ satisfies —B—^2 ¡ D_2(D_2+2), then $f$ is the composition of a Hopf fibration with a Möbius transformation $g:\mathbb{S}^{2}\to\mathbb{S}^{2}$ (up to conjugation).
Remark 1.1.
Seeking for a close analogy with the case of minimal hypersurfaces, one may presume that the Hopf fibration plays a role similar to that of the Clifford torus. However, this doesn’t seem to be the case for Theorem A, as the Hopf fibration satisfies (1.1) with strict inequality for any $a>0$, being $|B|^{2}=16$ and $\operatorname{D}_{2}=4$.
For harmonic maps with totally geodesic fibers, this result can be improved to:
Theorem B.
Suppose that $f:\mathbb{S}^{3}\to\mathbb{S}^{2}$ is a harmonic map with totally geodesic fibers. If the squared norm of its Hessian $B$ satisfies
$|B|^{2}\leq 2\operatorname{D}_{2}(\operatorname{D}_{2}+2),$
then $f$ is the composition of a Hopf fibration with a holomorphic map $g:\mathbb{S}^{2}\to\mathbb{S}^{2}$ (up to conjugation).
Remark 1.2.
From a technical viewpoint, there is a significant difference with the case of minimal hypersurfaces. If $f:M^{n}\to\mathbb{S}^{n+1}$ is a minimal hypersurface whose second fundamental form $B$ satisfies $|B|^{2}\leq n$, the rigidity result in [simons, lawson, CdCK] uses the maximum principle applied to Simons’ formula
$\frac{1}{2}\Delta|B|^{2}=|\nabla B|^{2}+n|B|^{2}-|B|^{4}.$
The Clifford torus is obtained in the case $|B|^{2}=n$ by looking at minimal hypersurfaces satisfying $|\nabla B|^{2}=0$. In our setting, for harmonic maps $f:\mathbb{S}^{3}\to\mathbb{S}^{2}$, one can compute the Laplacian of $|B|^{2}$ and obtain a more complicated formula of the type
$\Delta|B|^{2}=2|\nabla B|^{2}\ +\ (\text{terms of order zero in $B$}).$
However, if $f$ is the Hopf fibration then $|\nabla B|^{2}=96$, hence discarding the term one loses much information. Furthermore, the refined Kato inequality for harmonic maps
$|\nabla B|^{2}\geq\frac{3}{2}\big|\nabla|B|\big|^{2}$
is of no help, since for the Hopf fibration the right-hand side is zero. For these reasons, our pinching Theorems A and B will instead be obtained by applying maximum principles to carefully chosen combinations of the singular values of $f$, without differentiating $B$.
In our last main result, we consider harmonic maps whose singular values satisfy a suitable identity, without imposing bounds on $B$. In particular, the theorem applies to characterize harmonic maps $f:\mathbb{S}^{3}\to\mathbb{S}^{2}$ with constant energy density $|df|^{2}$ or constant 2-dilation $\operatorname{D}_{2}$. To state the result, recall that a function $f:C\subseteq\mathbb{R}^{n}\to\mathbb{R}$ is said to be $C^{1}$ on a subset $C$ if it can be extended to a $C^{1}$-function on an open set containing $C$.
Theorem C.
Suppose that $f:\mathbb{S}^{3}\to\mathbb{S}^{2}$ is a non-constant harmonic map whose two largest singular values $\lambda_{1}\geq\lambda_{2}$ satisfy ${W}(\lambda_{1},\lambda_{2})=0,$ for some function $W$ which is continuous on the set
$V=\big\{(x,y)\in\mathbb{R}^{2}\ :\ x\geq y\geq 0\},$
of class $C^{1}$ on $V\cap\{x>y\}\cap W^{-1}(0)$ and there it satisfies
$y\frac{\partial{W}}{\partial x}+x\frac{\partial{W}}{\partial y}\neq 0,\quad\text{for all}\,\,\,(x,y)\in V\cap\{x>y\}\cap W^{-1}(0)$ (1.2)
Then, $f$ is the composition of a Hopf fibration with a holomorphic map $g:\mathbb{S}^{2}\to\mathbb{S}^{2}$ (up to conjugation).
Theorem D.
Suppose that $f:\mathbb{S}^{3}\to\mathbb{S}^{2}$ is a non-constant harmonic map with constant energy density or constant $2$-dilation. Then, $f$ is a Hopf fibration.
Let us conclude by pointing out some additional characterizations of the Hopf fibration within certain classes of maps from $\mathbb{S}^{3}$ to $\mathbb{S}^{2}$. We highlight the following three:
  • Due to a result by Escobales [escobales], the (scaled) Hopf fibration is the only Riemannian submersion from $\mathbb{S}^{3}$ to $\mathbb{S}^{2}(1/2)$ up to isometry.
  • DeTurck, Gluck and Storm [Turck] proved that a map $f:\mathbb{S}^{3}\to\mathbb{S}^{2}$ with non-zero Hopf invariant and singular values less than or equal to $2$ must coincide with a Hopf fibration. In particular, up to isometries of domain and range, the Hopf fibration is the unique Lipschitz constant minimizer in its homotopy class.
  • The Hopf fibration has $2$-dilation $\operatorname{D}_{2}=4$. Recently, Guth and Lee [lee] proved that the Hopf invariant of a map $f:\mathbb{S}^{3}\to\mathbb{S}^{2}$ satisfies
    $|{\rm Hopf}(f)|\leq\frac{1}{16}\max_{\mathbb{S}^{3}}\operatorname{D}_{2}^{2}.$
    Consequently, if $\max_{\mathbb{S}^{3}}\operatorname{D}_{2}<4$ then $f$ is homotopically trivial.
Acknowledgements. L.M. is supported by the PRIN project no. 20225J97H5 “Differential-geometric aspects of manifolds via Global Analysis”.
Conflict of Interests. The authors have no conflict of interest.
Data availability statement. No data was generated by this research.
2. Basics on maps from $\mathbb{S}^{3}$ to $\mathbb{S}^{2}$
Let $f:U\subset\mathbb{S}^{3}\to\mathbb{S}^{2}$ be a smooth map, $U$ being an open subset of $\mathbb{S}^{3}$. The Euclidean metric on $\mathbb{R}^{4}$ is denoted by $\langle\cdot\,$$\cdot\rangle,$ and the same notation is used for the induced metrics on $\mathbb{S}^{3}$ and $\mathbb{S}^{2}$. Let $D$ be the Levi-Civita connection of $\mathbb{R}^{4}$, and $\nabla$ the Levi-Civita connections of $\mathbb{S}^{3}$ and $\mathbb{S}^{2}$. We also use the same symbol $\nabla$ for the pullback connection on the bundle $f^{*}T\mathbb{S}^{2}$. Denote by
$\mathcal{V}=\ker df\quad\text{and}\quad\mathcal{H}=\mathcal{V}^{\perp},$
respectively, the vertical and horizontal distributions.
2.1. The singular value decomposition
Denote by
$\lambda_{1}\geq\lambda_{2}\geq\lambda_{3}$
the singular values of $df$, i.e. the eigenvalues of $(df^{*}df)^{1/2}$, where $df^{*}$ is the transpose of $df$ with respect to the given metrics. Note that the singular values are continuous functions and that $\lambda_{3}\equiv 0$ for dimensional reasons. Suppose at the moment that on $U$ the kernel $\mathcal{V}$ of $df$ is a line bundle. Hence, in $U$ the two largest singular values of $df$ are non-zero. For simplicity, let us denote them by
$\lambda\doteq\lambda_{1}\geq\lambda_{2}\doteq\mu>0.$
At points where
$\lambda>\mu>0$
the singular values are even smooth, as the corresponding eigendistributions of $df^{*}df$ locally have constant dimensions, see [nomizu]. Fix $p\in U$ and consider an orthonormal basis $\{\alpha_{1},\alpha_{2},\alpha_{3}\}$ of $T_{p}\mathbb{S}^{3}$ and an orthonormal basis $\{\beta_{4},\beta_{5}\}$ of $T_{f(p)}\mathbb{S}^{2}$ such that
$\begin{array}[]{ccc}df(\alpha_{1})=\lambda\beta_{4},&df(\alpha_{2})=\mu\beta_{5},&df(\alpha_{3})=0.\end{array}$ (2.1)
The map $f:U\subset\mathbb{S}^{3}\to\mathbb{S}^{2}$ is called:
  1. (1)
    Horizontally conformal if $df:\mathcal{H}\to T\mathbb{S}^{2}$ is conformal or, equivalently, if and only if its singular values satisfy $\lambda\equiv\mu>0$ everywhere on $U$.
  2. (2)
    Weakly conformal if its two largest singular values satisfy $\lambda\equiv\mu\geq 0$, everywhere on $U$.
Note that in the case of weakly conformal maps we allow points where the differential of $f$ vanishes completely.
There are several quantities which encode information about how far $f:U\subset\mathbb{S}^{3}\to\mathbb{S}^{2}$ is from being weakly conformal. Consider the squared norm $u$ of the differential of $f$. In terms of the singular values of $f$ the function $u$ has the form
$u\doteq|df|^{2}=\lambda^{2}+\mu^{2}.$ (2.2)
The function $u$ is also known in the literature as (twice) the energy density of $f$. Since $\mathcal{H}$ is 2-dimensional, it possesses a complex structure $J_{\mathcal{H}}$. Let $\omega_{\mathbb{S}^{2}}$ be the volume form of $\mathbb{S}^{2}$. One can readily check that the $2$-form $f^{*}\omega_{\mathbb{S}^{2}}$ is non-zero only on $\mathcal{H}$. Consequently, it must be a multiple of the volume form $\omega_{\mathcal{H}}$ of $\mathcal{H}$. This means that there exists a smooth function $v$ on $U$ such that
$f^{*}\omega_{\mathbb{S}^{2}}\doteq v\,\omega_{\mathcal{H}}.$ (2.3)
In terms of the singular values of $f$ we have that
$|v|=\operatorname{D}_{2}=\lambda\mu,$ (2.4)
whence $v$ extends continuously to the entire $\mathbb{S}^{3}$ by setting $v=0$ at points where the dimension of $\mathcal{H}$ is less than 2. Without loss of generality we may assume that $v$ is positive on $U$. Observe that
$w\doteq u-2|v|=(\lambda-\mu)^{2}$
is a non-negative function. In particular, if at some point $p_{0}\in U$ it holds that $w(p_{0})=0,$ then $\lambda(p_{0})=\mu(p_{0}).$ Hence $w$ measures how much $f$ deviates from being weakly conformal.
2.2. The Hessian and the tension field
The Hessian of $f:\mathbb{S}^{3}\to\mathbb{S}^{2}$ is defined to be the symmetric vector valued tensor $B$ given by
$B(X,Y)\doteq\nabla_{X}df(Y)-df(\nabla_{X}Y),\quad\text{for}\,\,\,X,Y\in\mathfrak{X}(\mathbb{S}^{3}),$
where the first $\nabla$ stands for the natural connection on the pull-back bundle $f^{*}T\mathbb{S}^{2}$. By straightforward computations, it follows that
$\begin{array}[]{lcl}\displaystyle(\nabla_{X}B)(Y,Z)-(\nabla_{Z}B)(X,Y)&=&\displaystyle\big\{\langle X,Y\rangle-\langle df(X),df(Y)\rangle\big\}df(Z)\\[8.5359pt] &&\displaystyle-\big\{\langle Y,Z\rangle-\langle df(Y),df(Z)\rangle\big\}df(X),\end{array}$ (2.5)
for all $X,Y,Z\in\mathfrak{X}(\mathbb{S}^{3})$. Equation (2.5) will be referred to as the Codazzi equation. The trace of $B$ is called the tension field $\tau(f)$ of $f$. If $\tau(f)$ is identically zero the map $f$ is called harmonic. It is well-known that a harmonic map is real-analytic. Throughout the paper we adopt the following terminology:
$B_{ij}=B(\alpha_{i},\alpha_{j}),\qquad b_{ij}^{c}\doteq B^{c}(\alpha_{i},\alpha_{j})\doteq\langle B(\alpha_{i},\alpha_{j}),\beta_{c}\rangle,$
where $i,j,k\in\{1,2,3\}$ and $c\in\{4,5\}$. Moreover, at points where the frame $\{\alpha_{1},\alpha_{2},\alpha_{3}\}$ is smooth we set
$\omega_{ij}(\alpha_{k})\doteq\langle\nabla_{\alpha_{k}}\alpha_{i},\alpha_{j}\rangle.$
3. The Hopf fibration and harmonic unit vector fields
3.1. The Hopf fibration
Let us briefly describe the Hopf fibration, following the exposition of Pinkall [pinkall]. Identify $\mathbb{R}^{4}$ with the quaternions
$\mathbb{H}\doteq\{x_{0}+x_{1}i+x_{2}j+x_{3}k:x_{0},x_{1},x_{2},x_{3}\in\mathbb{R}\},$
which form an associative algebra with $1$ as the multiplicative unit via
$i^{2}=j^{2}=k^{2}=ijk=-1.$
Denote by $\operatorname{Re}\mathbb{H}$ the $1$-dimensional linear subspace spanned by the element $1$, and by $\operatorname{Im}\mathbb{H}$ the orthogonal complement of $\operatorname{Re}\mathbb{H}$. Given an element
$x=x_{0}+x_{1}i+x_{2}j+x_{3}k$
we define its conjugate as the element
$\overline{x}\doteq x_{0}-x_{1}i-x_{2}j-x_{3}k.$
Multiplication by $i$, $j$ and $k$ gives rise to complex structures $J_{1}$, $J_{2}$ and $J_{3}$ on $\mathbb{H}$ by
$J_{1}x\doteq i\cdot x,\quad J_{2}x\doteq j\cdot x\quad\text{and}\quad J_{3}x\doteq k\cdot x,\quad\text{for all}\,\,\,x\in\mathbb{H}.$
Consider the unit sphere $\mathbb{S}^{3}\subset\mathbb{H}$ as the subset of quaternions of length $1$, and identify $\mathbb{S}^{2}$ with the unit sphere in the subspace of $\mathbb{H}$ spanned by $1$, $j$ and $k$. Let $\tilde{\cdot}$ denote the anti-automorphism that fixes $1$, $j$ and $k$ and sends $i$ to i-i, i.e.
$x_{0}+x_{1}i+x_{2}j+x_{3}k=x\mapsto\tilde{x}=x_{0}-x_{1}i+x_{2}j+x_{3}k.$
Lemma 3.1.
Let $\pi:\mathbb{S}^{3}\to\mathbb{H}$ be the smooth map given by $\pi(p)\doteq\tilde{p}\cdot p$, for each $p\in\mathbb{S}^{3}.$ Then:
  1. (1)
    If $(x_{0},x_{1},x_{2},x_{3})\in\mathbb{S}^{3}\subset\mathbb{H}$, the map $\pi$ has the form π(x_0,x_1,x_2,x_3)=(x_0^2+x_1^2-x_2^2-x_3^2,2(x_0x_2+x_1x_3),2(x_0x_3-x_1x_2)).
  2. (2)
    The map $\pi:\mathbb{S}^{3}\to\mathbb{S}^{2}$ is a surjective submersion.
  3. (3)
    $\pi((\cos\theta+i\sin\theta)p)=\pi(p)$, for $\theta\in[0,2\pi]$ and $p\in\mathbb{S}^{3}$.
  4. (4)
    $\{\alpha_{1}(p)=j\cdot p,\alpha_{2}(p)=k\cdot p,\alpha_{3}=i\cdot p\}$ is an orthonormal frame on $\mathbb{S}^{3}$.
  5. (5)
    $\mathcal{V}={\rm ker}\,d\pi=\operatorname{span}\{\alpha_{3}\}$ and $\mathcal{H}=\mathcal{V}^{\perp}=\operatorname{span}\{\alpha_{1},\alpha_{2}\}.$ In particular, the integral curves of $\alpha_{3}$ are great circles.
  6. (6)
    The singular values of $\pi$ are constant; namely $\lambda_{1}=2=\lambda_{2}$ and $\lambda_{3}=0$.
  7. (7)
    The map $f\doteq(1/2)\pi:\mathbb{S}^{3}\to\mathbb{S}^{2}(1/2)$ is a Riemannian submersion.
The map $f$ described in the lemma above is known as the standard Hopf fibration and the associated vector field $\alpha_{3}=i\cdot p$ as the standard Hopf vector field. Any composition of $f$ with isometries or dilations of the corresponding spheres is called a Hopf fibration. Furthermore, any unit vector field $\zeta$ on $\mathbb{S}^{3}$ that can be expressed in the form
$\mathbb{S}^{3}\ni p\mapsto\zeta_{p}=J(p)\in T_{p}\mathbb{S}^{3},$
where $J$ is a complex structure on $\mathbb{R}^{4}$, is referred to in the literature as a Hopf vector field. Observe that the Hopf fibration is precisely the quotient map associated with the great circle fibration generated by the corresponding Hopf vector field.
In the next lemma we collect some further properties of a Hopf fibration.
Lemma 3.2.
Let $\pi:\mathbb{S}^{3}\to\mathbb{S}^{2}$ be the map given in Lemma 3.1 and $\{\alpha_{1},\alpha_{2},\alpha_{3}=\zeta\}$ the globally defined orthonormal frame on $\mathbb{S}^{3}$ arising from the singular value decomposition of $\pi$. Then:
  1. (1)
    The $(1,1)$-tensor $\varphi$ given by φ(X)-_Xζ is a complex structure on $\mathcal{H}$. Moreover, $\zeta$ satisfies the differential equation
    $\Delta\zeta+|\nabla\zeta|^{2}\zeta=0,$
    where $\Delta$ is the rough Laplacian on vector fields.
  2. (2)
    The map $\pi$ is harmonic, and the components of its Hessian with respect to the bases $\{\alpha_{1},\alpha_{2},\alpha_{3};\beta_{4},\beta_{5}\}$ arising from the singular value decomposition of $\pi$ are given by B^4= ( 00000-20-20 )  and B^5= ( 002000200 ), respectively. Moreover —B—^2=16 and —B—^2=6—B—^2=96.
  3. (3)
    Suppose that $g:\mathbb{S}^{2}\to\mathbb{S}^{2}$ is a conformal map. Then the composition $h=g\circ\pi$ is again a harmonic map with totally geodesic fibers. Moreover, the Hessian $B_{h}$ of $h$ is given by B^4_h=2 ( 2β_4(σ)2β_5(σ)02β_5(σ)-2β_4(σ)σ0σ0 )  and B^5_h=2 ( -2β_5(σ)2β_4(σ)2β_4(σ)2β_5(σ)000 ), where $\sigma$ is the (double) singular value of $g$. Moreover,
    $|B_{h}|^{2}=16(\sigma^{2}+4|\nabla\sigma|^{2}).$ (3.1)
    If $g$ is neither constant nor an isometry, then
    $\max|B_{h}|^{2}>16.$ (3.2)
Proof.
(1) Since the integral curves of $\zeta$ are great circles it follows that
$\varphi(\zeta)=0.$
Moreover, from the Weingarten formula, we have that
$\varphi(\alpha_{1})=-\nabla_{\alpha_{1}}\zeta=-D_{\alpha_{1}}(J_{1}x)=-J_{1}D_{\alpha_{1}}x=-J_{1}\alpha_{1}=-J_{1}J_{2}x=-J_{3}x=-\alpha_{2}.$
Similarly,
$\varphi(\alpha_{2})=\alpha_{1}.$
The formula for the Laplacian of $\zeta$ follows by straightforward computations.
(2) By the Koszul formula and the fact that $\pi$ has constant singular values we deduce that
$B(X,Y)=\nabla_{X}d\pi(Y)-d\pi(\nabla_{X}Y)=0,\quad\text{for all}\,\,X,Y\in\Gamma(\mathcal{H})$ (3.3)
for details see also [baird1, Lemma 4.5.1, page 119]. Hence, $B$ vanishes on the horizontal bundle. Since integral curves of $\zeta$ are geodesics and it belongs to the kernel of $d\pi$, we obtain that
$B(\zeta,\zeta)=0.$ (3.4)
and
$B(\zeta,X)=-d\pi(\nabla_{X}\zeta)=d\pi(\varphi(X)),\quad\text{for any}\,\,X\in\Gamma(\mathcal{H}).$ (3.5)
Differentiating (3.3), (3.4) and (3.5) we get the estimates for $|B|$ and $|\nabla B|$.
(3) Let us denote by $\sigma$ the conformal factor of $g$. By a straightforward computation, we get
$B_{h}(X,Y)=B_{g}(d\pi(X),d\pi(Y))+dg(B(X,Y)),$ (3.6)
for any vector fields $X$ and $Y$ of the sphere $\mathbb{S}^{3}$. Hence,
$B_{h}(\zeta,\zeta)=B_{g}(d\pi(\zeta),d\pi(\zeta))+dg(B(\zeta,\zeta))=0.$
Consider at $h(x)$ an orthonormal frame $\{{\tilde{\beta}_{4}},{\tilde{\beta}}_{5}\}$ such that
$dg(\beta_{4})=\sigma{\tilde{\beta}}_{4}\quad\text{and}\quad dg(\beta_{5})=\sigma{\tilde{\beta}}_{5}.$
Again by Koszul’s formula we get, away from the zero set of $\sigma$, that
$B_{g}(X,Y)=X(\log\sigma)dg(Y)+Y(\log\sigma)dg(X)-\langle X,Y\rangle dg(\nabla\log\sigma),$
where $X$ and $Y$ are tangent vectors of $\mathbb{S}^{2}$. Hence, from (3.6), (3.3), (3.4) and (3.5) we deduce that
$B_{g}(\alpha_{3},\alpha_{3})=0,\quad B_{g}(\alpha_{1},\alpha_{3})=2\sigma{\tilde{\beta}}_{5}\quad\text{and}\quad B_{g}(\alpha_{2},\alpha_{3})=-2\sigma{\tilde{\beta}}_{4}.$
Furthermore
$B_{h}(\alpha_{1},\alpha_{1})=4\beta_{4}(\sigma)\tilde{\beta}_{4}-4\beta_{5}(\sigma){\tilde{\beta}}_{5}=-B_{h}(\alpha_{2},\alpha_{2})$
and
$B_{h}(\alpha_{1},\alpha_{2})=4\beta_{5}(\sigma)\tilde{\beta}_{4}+4\beta_{4}(\sigma){\tilde{\beta}}_{5}.$
Assume now that $g$ is neither constant nor an isometry. We claim that
$\min\sigma^{2}<1<\max\sigma^{2}.$
Indeed, suppose to the contrary that $\max\sigma^{2}\leq 1$, i.e. $g$ is a weakly length decreasing map. Because $g$ is conformal, its graph is a minimal submanifold of $\mathbb{S}^{2}\times\mathbb{S}^{2}$; see for example [eells]. According to [savas4, Theorem A], $g$ must be constant or an isometry, a contradiction. Assume now that $\min\sigma^{2}\geq 1$. Then $g$ is a local diffeomorphism. Since $\mathbb{S}^{2}$ is simply connected, $g$ is a global diffeomorphism. Then $g^{-1}$ is a weakly length decreasing minimal diffeomorphism; see for example [schoen2]. Again due to [savas4, Theorem A], $g$ must be an isometry. Since $\nabla\sigma$ vanishes at maximum points of $\sigma$, it follows from (3.1) that $|B_{h}|^{2}$ takes values above $16$. This completes the proof. ∎
3.2. Harmonic unit vector fields on $\mathbb{S}^{3}$
In the previous section we proved that the Hopf vector field $\zeta$ satisfies the equation
$\Delta\zeta+|\nabla\zeta|^{2}\zeta=0,$ (3.7)
where $\Delta$ is the rough Laplacian of $\zeta$. It turns out that (3.7) is the Euler-Lagrange equation of the energy functional
$E(\zeta)\doteq\frac{3}{2}{\rm vol}(\mathbb{S}^{3})+\frac{1}{2}\int_{\mathbb{S}^{3}}|\nabla\zeta|^{2}d\mu$
in the space of unit vector fields on $\mathbb{S}^{3}$, where $d\mu$ denotes the volume element and ${\rm vol}(\mathbb{S}^{3})$ is the volume of $\mathbb{S}^{3}$; see [wiegmink, wood]. Critical points of $E$ for variations by unit vector fields are called harmonic unit vector fields. It is conjectured that the Hopf vector fields are the only harmonic unit vector fields. This conjecture remains open in general, but Fourtzis, Markellos, and Savas-Halilaj [savas2, Theorem B & C] proved it under the additional assumption that $\zeta$ has totally geodesic integral leaves. To state the result, we first recall the following topological lemma.
Lemma 3.3.
Let $\zeta$ be a unit vector field on $\mathbb{S}^{3}$ with compact integral leaves. Then, the quotient space obtained by identifying points on the same leaves of $\zeta$ is a smooth compact surface diffeomorphic to $\mathbb{S}^{2}$, and the quotient map $\pi:\mathbb{S}^{3}\to\mathbb{S}^{2}$ is a smooth fiber bundle (in particular, a submersion).
Proof.
Let us denote by $\mathbb{M}^{2}$ the space of leaves of the foliation generated by the integral curves of $\zeta$, and denote by $\pi:\mathbb{S}^{3}\to\mathbb{M}^{2}$ the corresponding quotient map, i.e. the map which identifies points of the same integral curve of $\zeta$. Note that $\mathbb{M}^{2}=\pi(\mathbb{S}^{3})$ with the quotient topology is compact and connected. Since $\zeta$ has compact leaves, the distribution it generates is regular in the sense of [palais, Chapter 1], thus by [palais, Corollary 5 p.24] the triple
$\mathbb{S}^{1}\hookrightarrow\mathbb{S}^{3}\overset{\pi}{\longrightarrow}\mathbb{M}^{2}$
is $\mathbb{S}^{1}$ fiber bundle and $\mathbb{M}^{2}$ is a smooth $2$-dimensional manifold. Moreover, by the long exact homotopy sequence of a fiber bundle (see [hatcher, Theorem 4.41 and Proposition 4.48])
$\ldots\to\pi_{1}(\mathbb{S}^{3})\to\pi_{1}(\mathbb{M}^{2})\to\pi_{0}(\mathbb{S}^{1})\to\pi_{0}(\mathbb{S}^{3})\to 0$
we deduce that $\mathbb{M}^{2}$ is simply connected. Therefore, $\mathbb{M}^{2}$ is diffeomorphic to the sphere $\mathbb{S}^{2}$. ∎
Theorem 3.4 ([savas2]).
Let $\zeta\in\mathfrak{X}(\mathbb{S}^{3})$ be a unit vector field whose integral curves are great circles and let $\pi:\mathbb{S}^{3}\to\mathbb{S}^{2}$ be the corresponding quotient map. Equip the spheres with their standard round metrics. Then:
  1. (1)
    The map $\pi$ is harmonic if and only if $\zeta$ is a harmonic unit vector field.
  2. (2)
    If $\zeta$ is globally defined, then it coincides with a Hopf vector field and the corresponding quotient map $\pi:\mathbb{S}^{3}\to\mathbb{S}^{2}$ is a Hopf fibration.
4. Bochner-Weitzenböck formulas
In the sequel we will compute the gradients and Laplacians of the functions $u$, $v$ and $w$ defined in the previous section, in the case where $f$ is a harmonic map.
Lemma 4.1.
Let $f:U\subset\mathbb{S}^{3}\to\mathbb{S}^{2}$ be a submersion and $k\in\{1,2,3\}$. On the open and dense subset of $U$ where $\lambda$ and $\mu$ are smooth we have that
$\alpha_{k}(\lambda)=b_{1k}^{4}\quad\text{and}\quad\alpha_{k}(\mu)=b_{2k}^{5}.$
Moreover,
$\lambda\omega_{13}(\alpha_{k})=b_{3k}^{4}\quad\text{and}\quad\mu\omega_{23}(\alpha_{k})=b_{3k}^{5}$
and
$(\lambda^{2}-\mu^{2})\omega_{12}(\alpha_{k})=\lambda b_{2k}^{4}+\mu b_{1k}^{5},$
where the indices are with respect to the frames arising from the singular value decomposition.
Proof.
Differentiating the equation $\lambda^{2}=|df(\alpha_{1})|^{2}$ with respect to $\alpha_{k}$ we get
$\lambda\alpha_{k}(\lambda)=\langle B(\alpha_{k},\alpha_{1})+df(\nabla_{\alpha_{k}}\alpha_{1}),df(\alpha_{1})\rangle=\langle B(\alpha_{k},\alpha_{1}),df(\alpha_{1})\rangle=\lambda\,b^{4}_{1k}.$
In a similar manner, by differentiating the equation $\mu^{2}=|df(\alpha_{2})|^{2},$ we obtain that
$\alpha_{k}(\mu)=b_{2k}^{5}.$
On the other hand, because $\alpha_{3}$ is in the kernel of $df$, we get that
$B(\alpha_{k},\alpha_{3})=-df(\nabla_{\alpha_{k}}\alpha_{3})=\lambda\omega_{13}(\alpha_{k})\beta_{4}+\mu\omega_{23}(\alpha_{k})\beta_{5}.$
Differentiating
$\langle df(\alpha_{1}),df(\alpha_{2})\rangle=0$
with respect to $\alpha_{k}$ we get the expression for $\omega_{12}(\alpha_{k})$. This concludes the proof. ∎
Lemma 4.2.
Let $f:U\subset\mathbb{S}^{3}\to\mathbb{S}^{2}$ be a harmonic map. On the open set where $\lambda$ is different from $\mu$, we have that $\lambda^{2}$ and $\mu^{2}$ are smooth functions, and their Laplacians are given by
$\frac{1}{2}\Delta(\lambda^{2})=(2-\mu^{2})\lambda^{2}+\sum_{k=1}^{3}\frac{(\lambda b_{2k}^{4}+\mu b_{1k}^{5})^{2}}{\lambda^{2}-\mu^{2}}+\sum_{k=1}^{3}|B_{1k}|^{2}+\sum_{k=1}^{3}(b_{3k}^{4})^{2},$
and
$\frac{1}{2}\Delta(\mu^{2})=(2-\lambda^{2})\mu^{2}-\sum_{k=1}^{3}\frac{(\lambda b_{2k}^{4}+\mu b_{1k}^{5})^{2}}{\lambda^{2}-\mu^{2}}+\sum_{k=1}^{3}|B_{2k}|^{2}+\sum_{k=1}^{3}(b_{3k}^{5})^{2},$
where the indices are with respect to the frames arising from the singular value decomposition.
Proof.
On the open set $V$ where $\lambda>\mu$ the multiplicity of $\lambda^{2}$ does not change. Thus $\lambda^{2}$ is smooth [nomizu]. Since $u$ is smooth on $\mathbb{S}^{3}$ it follows that $\mu^{2}$ is smooth on $V$ as well. We have
$\frac{1}{2}\Delta(\lambda^{2})$ $=\frac{1}{2}\sum_{k=1}^{3}(\nabla_{\alpha_{k}}\nabla_{\alpha_{k}}-\nabla_{\nabla_{\alpha_{k}}\alpha_{k}})\langle df(\alpha_{1}),df(\alpha_{1})\rangle$
$=\sum_{k=1}^{3}\langle\nabla_{\alpha_{k}}B(\alpha_{1},\alpha_{k}),df(\alpha_{1})\rangle+\sum_{k=1}^{3}\langle B(\alpha_{1},\alpha_{k}),B(\alpha_{1},\alpha_{k})+df(\nabla_{\alpha_{k}}\alpha_{1})\rangle$
$\quad-\sum_{k=1}^{3}\langle B(\alpha_{1},\nabla_{\alpha_{k}}\alpha_{k}),df(\alpha_{1})\rangle$
$=\sum_{k=1}^{3}\langle(\nabla_{\alpha_{k}}B)(\alpha_{1},\alpha_{k})+B(\nabla_{\alpha_{k}}\alpha_{1},\alpha_{k}),df(\alpha_{1})\rangle+\sum_{k=1}^{3}\big(|B_{1k}|^{2}+\omega_{12}(\alpha_{k})\mu b_{1k}^{5}\big).$
By the Codazzi equation (2.5) and Lemma 4.1 we get that
$\frac{1}{2}\Delta(\lambda^{2})$ $=(2-\mu^{2})\lambda^{2}+\sum_{k=1}^{3}\frac{(\lambda b_{2k}^{4}+\mu b_{1k}^{5})^{2}}{\lambda^{2}-\mu^{2}}+\sum_{k=1}^{3}|B_{1k}|^{2}+\sum_{k=1}^{3}(b_{3k}^{4})^{2}.$
In a similar manner
$\frac{1}{2}\Delta(\mu^{2})$ $=\sum_{k=1}^{3}\langle\nabla_{\alpha_{k}}B(\alpha_{2},\alpha_{k}),df(\alpha_{2})\rangle+\sum_{k=1}^{3}\langle B(\alpha_{2},\alpha_{k}),B(\alpha_{2},\alpha_{k})+df(\nabla_{\alpha_{k}}\alpha_{2})\rangle$
$\quad-\sum_{k=1}^{3}\langle B(\alpha_{2},\nabla_{\alpha_{k}}\alpha_{k}),df(\alpha_{2})\rangle$
$=\sum_{k=1}^{3}\langle(\nabla_{\alpha_{K}}B)(\alpha_{2},\alpha_{k})+B(\nabla_{\alpha_{k}}\alpha_{2},\alpha_{k}),df(\alpha_{2})\rangle+\sum_{k=1}^{3}\big(|B_{2k}|^{2}+\omega_{21}(\alpha_{k})\lambda b_{2k}^{4})\big)$
$=(2-\lambda^{2})\mu^{2}-\sum_{k=1}^{3}\frac{(\lambda b_{2k}^{4}+\mu b_{1k}^{5})^{2}}{\lambda^{2}-\mu^{2}}+\sum_{k=1}^{3}|B_{2k}|^{2}+\sum_{k=1}^{3}(b_{3k}^{5})^{2},$
which completes the proof. ∎
Making use of Lemmata 4.1 and 4.2, we obtain the following by direct computations.
Lemma 4.3.
Let $f:U\subset\mathbb{S}^{3}\to\mathbb{S}^{2}$ be a harmonic map. The gradient and the Laplacian of the function $u$ defined in (2.2) are given by
$\nabla u=2\sum_{k=1}^{3}(\lambda b_{1k}^{4}+\mu b_{2k}^{5})\alpha_{k}$ (4.1)
where the indices are with respect to the frames of the singular value decomposition. Moreover
$\Delta u=2|B|^{2}+4(u-v^{2}),$
where $v$ is the function given in (2.4). In particular $v^{2}$ is a real analytic function on $U$.
Lemma 4.4.
Let $f:U\subset\mathbb{S}^{3}\to\mathbb{S}^{2}$ be a harmonic submersion. The gradient and Laplacian of the product $\lambda\mu$ satisfy
$\nabla(\lambda\mu)=\sum_{k=1}^{3}(\mu b_{1k}^{4}+\lambda b_{2k}^{5})\alpha_{k}$
and
$\Delta(\lambda\mu)=\lambda\mu(4-\lambda^{2}-\mu^{2})+2\sum_{k=1}^{3}(b_{1k}^{4}b_{2k}^{5}-b_{1k}^{5}b_{2k}^{4})+\frac{\mu}{\lambda}\sum_{k=1}^{3}(b_{3k}^{4})^{2}+\frac{\lambda}{\mu}\sum_{k=1}^{3}(b_{3k}^{5})^{2},$
where the indices are with respect to the frames arising from the singular value decomposition.
As an immediate consequence of Lemmata 4.3 and 4.4 we obtain the following result.
Lemma 4.5.
Let $f:U\subset\mathbb{S}^{3}\to\mathbb{S}^{2}$ be a harmonic submersion. Consider the function
$w=(\lambda-\mu)^{2}\geq 0.$
At points where $w$ is smooth, its gradient and Laplacian are given by
$\nabla w=2(\lambda-\mu)\sum_{k=1}^{3}(b_{1k}^{4}-b_{2k}^{5})\alpha_{k}$
and
$\frac{1}{2}\Delta w=w(\lambda\mu+2)+|B|^{2}-2\sum_{k=1}^{3}(b_{1k}^{4}b_{2k}^{5}-b_{1k}^{5}b_{2k}^{4})-\frac{\mu}{\lambda}\sum_{k=1}^{3}(b_{3k}^{4})^{2}-\frac{\lambda}{\mu}\sum_{k=1}^{3}(b_{3k}^{5})^{2},$
where the indices are with respect to the frames arising from the singular value decomposition.
Moreover, by a direct computation we obtain the following lemma.
Lemma 4.6.
Let $f:U\subset\mathbb{S}^{3}\to\mathbb{S}^{2}$ be a harmonic map. Consider the function
$\varrho\doteq\sqrt{u^{2}-4v^{2}}=\lambda^{2}-\mu^{2}\geq 0.$
At points where $\varrho$ is smooth, its Laplacian is given by
$\frac{1}{2}\Delta\varrho=2\varrho+2\varrho^{-1}\sum_{k=1}^{3}(\lambda b_{2k}^{4}+\mu b_{1k}^{5})^{2}+|B_{11}|^{2}-|B_{22}|^{2}+2(b_{13}^{4})^{2}-2(b_{23}^{5})^{2}+(b_{33}^{4})^{2}-(b_{33}^{5})^{2},$
where the indices are with respect to the frames arising from the singular value decomposition.
5. Weakly conformal harmonic maps from $\mathbb{S}^{3}$ to $\mathbb{S}^{2}$
In this section we consider weakly conformal harmonic maps $f:U\subset\mathbb{S}^{3}\to\mathbb{S}^{2}$. If $U$ is the entire $\mathbb{S}^{3}$, such maps were classified by Baird and Wood in [baird2, Theorem 5.1]. They showed the following:
Theorem 5.1.
Suppose that $f:\mathbb{S}^{3}\to\mathbb{S}^{2}$ is a weakly conformal harmonic map. Then $f$ factors as the composition of a Hopf fibration $\pi:\mathbb{S}^{3}\to\mathbb{S}^{2}$ with a conformal map $g:\mathbb{S}^{2}\to\mathbb{S}^{2}$. In particular, if $f$ is a submersion, then $g$ is a Möbius transformation (up to conjugation).
For the sake of completeness, we include a proof of this result which combines Theorem 3.4 with a detailed analysis of the local structure of the locus where the differential of the map $f$ vanishes identically.
We first recall a couple of useful facts about weakly conformal submersions. Assume that $U\subset\mathbb{S}^{3}$ is an open subset and that $f:U\to f(U)\subset\mathbb{S}^{2}$ is a submersion. Then there exists a unique unit vector field $\zeta\in\mathfrak{X}(U)$ spanning $\ker(df)$ and such that $f^{*}\omega_{\mathbb{S}^{2}}\wedge\zeta_{\flat}$ is in the orientation of $\mathbb{S}^{3}$.
The following lemma is well-known in the literature, see for example [baird1, Proposition 4.5.3], but for the sake of completeness we include a short proof.
Lemma 5.2.
Let $f:U\subset\mathbb{S}^{3}\to\mathbb{S}^{2}$ be a weakly conformal submersion, defined in an open neighbourhood $U$ of $\mathbb{S}^{3}$. Then $f$ is harmonic if and only if its fibers are totally geodesic.
Proof.
We define $\zeta$ as above on $U$, and denote by $\lambda$ the non-zero (double) singular value of $df$. By Koszul’s formula we have that
$B(X,Y)=X(\log\lambda)df(Y)+Y(\log\lambda)df(X)-\langle X,Y\rangle df(\nabla\log\lambda),$ (5.1)
for any pair of horizontal vector fields $X,Y\in\Gamma(\mathcal{H})$. Furthermore,
$B(\zeta,\zeta)=\nabla_{\zeta}df(\zeta)-df(\nabla_{\zeta}\zeta)=-df(\nabla_{\zeta}\zeta)=-df(H_{\zeta}),$ (5.2)
where $H_{\zeta}$ is the geodesic curvature of the integral curves of $\zeta$. From (5.1) and (5.2) we readily see that
$\tau(f)=-df(H_{\zeta}).$
This completes the proof. ∎
We now show that $\zeta$ is in fact a harmonic unit vector field.
Lemma 5.3.
Let $f:U\subset\mathbb{S}^{3}\to\mathbb{S}^{2}$ be a weakly conformal harmonic submersion. Then the unit vector field $\zeta$ generating the vertical distribution $\mathcal{V}$ is harmonic.
Proof.
Consider a local orthonormal frame $\{\alpha_{1},\alpha_{2},\alpha_{3}\doteq\zeta\}$ and denote by $\lambda$ the non-zero (double) singular value of $df$. From Lemma 4.1 for $\lambda\equiv\mu$ we see that
$b_{1k}^{4}=b_{2k}^{5}\quad\text{and}\quad b_{1k}^{5}=-b_{2k}^{4},\quad\text{for}\,\,k\in\{1,2,3\}.$ (5.3)
Define the vector field
$Z\doteq\Delta\zeta+|\nabla\zeta|^{2}\zeta.$
We can easily see that $Z$ is perpendicular to $\zeta$. Indeed
$\langle Z,\alpha_{3}\rangle=\sum_{k=1}^{2}\langle\nabla_{\alpha_{k}}\nabla_{\alpha_{k}}\alpha_{3},\alpha_{3}\rangle+\sum_{k=1}^{2}|\nabla_{\alpha_{k}}\alpha_{3}|^{2}=-\sum_{k=1}^{2}\langle\nabla_{\alpha_{k}}\alpha_{3},\nabla_{\alpha_{k}}\alpha_{3}\rangle+\sum_{k=1}^{2}|\nabla_{\alpha_{k}}\alpha_{3}|^{2}=0.$
It remains to show that the horizontal component of $Z$ is zero as well, or equivalently that
$df(Z)=0.$
Using the Codazzi equation and the harmonicity of $f$, we obtain
$df(Z)$ $=\sum_{k=1}^{3}df\big\{\nabla_{\alpha_{k}}\nabla_{\alpha_{k}}\alpha_{3}-\nabla_{\nabla_{\alpha_{k}}\alpha_{k}}\alpha_{3}\big\}$
$=\sum_{k=1}^{3}\big\{\nabla_{\alpha_{k}}df(\nabla_{\alpha_{k}}\alpha_{3})-B(\alpha_{k},\nabla_{\alpha_{k}}\alpha_{3})-\nabla_{\nabla_{\alpha_{k}}\alpha_{k}}df(\alpha_{3})+B(\nabla_{\alpha_{k}}\alpha_{k},\alpha_{3})\big\}$
$=\sum_{k=1}^{3}\big\{\nabla_{\alpha_{k}}\nabla_{\alpha_{k}}df(\alpha_{3})-\nabla_{\alpha_{k}}B(\alpha_{k},\alpha_{3})-B(\alpha_{k},\nabla_{\alpha_{k}}\alpha_{3})+B(\nabla_{\alpha_{k}}\alpha_{k},\alpha_{3})\big\}.$
$=-\sum_{k=1}^{3}\big\{(\nabla_{\alpha_{k}}B)(\alpha_{3},\alpha_{k})-2B(\nabla_{\alpha_{k}}\alpha_{3},\alpha_{k})\big\}$
$=-\sum_{k=1}^{3}\big\{(\nabla_{\alpha_{3}}B)(\alpha_{k},\alpha_{k})+2\omega_{13}(\alpha_{k})B_{1k}+2\omega_{23}(\alpha_{k})B_{2k}\big\}$
$=-\sum_{k=1}^{3}\big\{2\omega_{13}(\alpha_{k})B_{1k}+2\omega_{23}(\alpha_{k})B_{2k}\big\}.$
By Lemma 4.1 for $\lambda\equiv\mu$ and (5.3) we deduce that
$df(Z)$ $=\frac{2}{\lambda}\sum_{k=1}^{3}(b^{4}_{k3}B_{1k}+b^{5}_{k3}B_{2k})$
$=\frac{2}{\lambda}(b_{13}^{4}b_{11}^{4}+b_{23}^{4}b_{12}^{4}+b_{13}^{5}b_{12}^{4}+b_{23}^{5}b_{22}^{4})\beta_{4}$
$\quad+\frac{2}{\lambda}(b_{13}^{4}b_{11}^{5}+b_{23}^{4}b_{12}^{5}+b_{13}^{5}b_{12}^{5}+b_{23}^{5}b_{22}^{5})\beta_{5}$
$=\frac{2}{\lambda}(b_{13}^{4}b_{11}^{4}+b_{23}^{4}b_{12}^{4}-b_{23}^{4}b_{12}^{4}-b_{13}^{4}b_{11}^{4})\beta_{4}$
$\quad+\frac{2}{\lambda}(b_{13}^{4}b_{11}^{5}+b_{23}^{4}b_{12}^{5}-b_{23}^{4}b_{12}^{5}-b_{13}^{4}b_{11}^{5})\beta_{5}$
$=0,$
which concludes the proof. ∎
We are ready to prove Theorem 5.1.
Proof of Theorem 5.1: Define
$U=\{p\in\mathbb{S}^{3}:\operatorname*{rank}(df_{p})=2\}.$
Note that $U$ is open and that $f:U\to f(U)\subset\mathbb{S}^{2}\subset\mathbb{R}^{3}$ is a weakly conformal submersion with totally geodesic fibers by Lemma 5.3. Let $\zeta$ be the unit vector field on $U$ defined as above. We denote the fibers of $f$ passing through $p$ by
$\mathcal{A}_{p}\doteq\{q\in\mathbb{S}^{3}:f(q)=f(p)\}.$
Since $f$ is real analytic, each $\mathcal{A}_{p}$ is the zero set of the real analytic function $h:U\to\mathbb{R}$ given by
$h(q)=\|f(q)-f(p)\|^{2}.$
By Łojasiewicz’s structure theorem [kr, Theorem 6.3.3], there exists a neighbourhood of $p$ where $\mathcal{A}_{p}$ decomposes as the disjoint union
$\mathcal{A}_{p}=\mathcal{V}_{0}\cup\mathcal{V}_{1}\cup\mathcal{V}_{2}\cup\mathcal{V}_{3},$
where $\mathcal{V}_{d}$, $0\leq d\leq 3$, is either empty or a finite disjoint union of $d$-dimensional real analytic sub-varieties without boundary. Moreover, up to shrinking the neighbourhood of $p$, if $\mathcal{V}_{d}\neq\emptyset$ then $\overline{\mathcal{V}}_{d}$ contains the union of all lower dimensional strata $\mathcal{V}_{j}$ for $0\leq j\leq d-1$. A point $q\in\mathcal{A}_{p}$ is called a regular point of dimension $d$ if there exists a neighborhood $\Omega$ of $q$ such that $\Omega\cap\mathcal{A}_{p}$ is a connected $d$-dimensional real analytic submanifold of $\Omega$. In particular, a regular point of dimension zero is an isolated point of $\mathcal{V}$. Otherwise, $q$ is called a singular point.
Remark 5.4.
By the last property we quote of Łojasiewicz’s theorem, each singular point is a limit of regular points lying in the same connected component of $\mathcal{A}_{p}$. In particular, $\mathcal{A}_{p}$ must have at least one regular point.
Since $|df|^{2}$ is a real analytic function as well, the above discussion also applies to
$\mathcal{A}_{df}\doteq\{p\in\mathbb{S}^{3}:|df_{p}|=0\}.$
Notice that
$df^{*}df\sim\begin{pmatrix}\lambda^{2}&0&0\\ 0&\lambda^{2}&0\\ 0&0&0\end{pmatrix}$,
therefore the rank of the differential can be either 0 or $2$ and $\mathcal{A}_{df}=\mathbb{S}^{3}\setminus U.$
Suppose henceforth that $f$ in non-constant.
Lemma 5.5.
The map $f$ is constant on connected components of $\mathcal{A}_{df}$.
Proof.
The result is trivial for isolated points, i.e. regular points of dimension 0. Let $p$ be a regular point of $\mathcal{A}_{df}$ of dimension $d\geq 1$. There exists an open connected neighborhood $\Omega$ of $p$ such that $\Omega\cap\mathcal{A}_{df}$ is a connected $d$-dimensional submanifold of $\Omega$. For any curve $\gamma:[0,1]\to\Omega\cap\mathcal{A}_{df}$ we have that $(f\circ\gamma)^{\prime}\equiv 0$ everywhere, therefore $f\circ\gamma$ is constant. Since any two points of $\Omega\cap\mathcal{A}_{df}$ can be connected by a smooth curve, it must be that $f$ is constant on connected sets of regular points of $\mathcal{A}_{df}$. This can be extended to include singular points since $f$ is continuous. ∎
Lemma 5.6.
The fibers of $f$ have no regular points of dimension $2$ or $3$.
Proof.
Regular points of dimension $3$ would force $f$ to be constant by analyticity. Let $p$ be a regular point of $\mathcal{A}_{p}$ of dimension $2$. Then there exists a connected open neighborhood $\Omega$ of $p$ such that $\mathcal{A}_{p}\cap\Omega$ is a $2$-dimensional submanifold of $\Omega$. Furthermore, notice that $\operatorname*{rank}df\leq 1$ implies $\operatorname*{rank}df=0$ on $\mathcal{A}_{p}\cap\Omega$. It follows that $f$ is a solution of the Cauchy problem
[…]
By the Cauchy-Kowalewsky theorem, this problem admits the unique solution $f=f(p)$ in $\Omega$. Therefore $f$ is again constant by analyticity, which is a contradiction. ∎
Lemma 5.7.
The fibers of $f$ consist of non-intersecting great circles in $\mathbb{S}^{3}$, and the vector field $\zeta$ admits a smooth extension on the entire $\mathbb{S}^{3}$.
Proof.
Concerning the extension, since $\zeta$ is smooth on $U$ we only have to consider points $p\in\mathcal{A}_{df}=\mathbb{S}^{3}\setminus U$. By Lemma 5.5 the connected component of $\mathcal{A}_{df}$ containing $p$ lies in $\mathcal{A}_{p}$.
Claim 1: $\mathcal{A}_{p}$ contains at least one great circle passing through $p$. In particular, $p$ cannot be a regular point of dimension 0 for $\mathcal{A}_{p}$.
[…]
Corollary to Claim 1: If $p$ is a regular point of dimension $1$ for $\mathcal{A}_{p}$, then any other sub-sequential limit $\overline{\gamma}_{\infty}:\mathbb{S}^{1}\to\mathbb{S}^{3}$ of $\gamma_{n}$ must satisfy
$\overline{\gamma}_{\infty}=\pm\gamma_{\infty}.$
Claim 2: If $p$ is a regular point of dimension $1$ for $\mathcal{A}_{p}$, there exists an open geodesic disc $D$ centered at $p$, transversal to the fibers of $f$, whose unit normal $\nu$ can be chosen such that $\nu|_{p}=\gamma^{\prime}_{\infty}(0)$.
Proof of Claim 2. Let
$\mathcal{H}_{p}=(T_{p}\gamma_{\infty}(\mathbb{S}^{1}))^{\perp}\subset T_{p}\mathbb{S}^{3}.$
Denote by $B_{0}^{\mathcal{H}}(r)$ the ball of radius $r$ in $\mathcal{H}_{p}$ centered at the origin. Suppose that none of the discs
$D_{n}\doteq\exp_{p}(B_{0}^{\mathcal{H}}(\tfrac{1}{n}))$
are transversal to the fibers of $f$. This implies that there exists a sequence of points $(p_{n})_{n\in\mathbb{N}}$ such that
[…]
Claim 3: If $p$ is a regular point of dimension $1$ for $\mathcal{A}_{p}$, then the entire sequence $(\gamma_{n})_{n\in\mathbb{N}}$ described in Claim 1 converges to the great circle $\gamma_{\infty}$.
Proof of Claim 3. Let $D$ be the disc given in Claim 2. We shrink $D$ if necessary, so that
$\langle\zeta,\nu\rangle\geq\delta>0,\quad\text{on $D\setminus\{p\}$}.$
The fact that this quantity does not approach zero near $p$ is guaranteed by the proof of Claim 2. We now extend $\nu$ by parallel transport along the geodesic fibers of $f$, so that
$\langle\zeta,\nu\rangle\geq\delta,\quad\text{on $\Omega\setminus\mathcal{A}_{p}$}.$
If $\overline{\gamma}_{\infty}:\mathbb{S}^{1}\to\mathbb{S}^{3}$ is the limit geodesic of another subsequence $(\gamma_{n_{l}})_{l\in\mathbb{N}}$, this implies that
$\langle\overline{\gamma}^{\prime}_{\infty}(0),\gamma^{\prime}_{\infty}(0)\rangle=\lim_{l\to\infty}\langle\zeta,\nu\rangle(p_{n_{l}})\geq\delta>0.$
It follows from the Corollary to Claim 1 that $\overline{\gamma}_{\infty}\equiv\gamma_{\infty},$ as claimed.
Claim 4: $\zeta$ admits a smooth extension to any $p\in\mathcal{A}_{df}$ which is a regular point of its fiber.
[…]
Claim 5: For each $p\in\mathbb{S}^{3}$, the set $\mathcal{A}_{p}$ is a disjoint union of great circles. In particular, $\mathcal{A}_{p}$ does not have singular points.
Proof of Claim 5. By Claims 1 and 3, we only need to consider the case where $p$ is a singular point of $\mathcal{A}_{p}$. By Łojasiewicz’s Theorem and Lemma 5.6, the point $p$ is in the closure of a union of disjoint curves. Each of them is therefore made up of regular points of $\mathcal{A}_{p}$ of dimension $1$ and thus, by Claim 1, it is a piece of great circle contained in $\mathcal{A}_{p}$. To prove the claim it is therefore enough to exclude the possibility that $\mathcal{A}_{p}$ contains two great circles $\sigma_{j}:\mathcal{S}^{1}\to\mathbb{S}^{3}$, $j\in\{1,2\}$, meeting transversely at $p$ at $t=0$, i.e.
$\sigma_{1}(0)=p=\sigma_{2}(0)\quad\text{and}\quad\langle\sigma^{\prime}_{1}(0),\sigma^{\prime}_{2}(0)\rangle\neq{\pm 1}.$
Note that $(f\circ\sigma_{1})\equiv(f\circ\sigma_{2})\equiv f(p).$ Let $q$ be a regular point of $\mathcal{A}_{p}$ in the image of $\sigma_{1}$. We have shown in Claim 4 that $\zeta$ admits a smooth extension in a neighborhood $\Omega$ of $q$. Let $D$ be a disk centered at $q$ as described in Claim 2, and proceed as in the proof of Claim 3 so that
$\langle\zeta,\nu\rangle\geq\delta>0,\quad\text{on $\Omega$}.$
In particular, $\zeta|_{D}$ is smooth, and $\zeta|_{\Omega}$ is obtained by parallelly transporting $\zeta|_{D}$ along the geodesic fibers of the map $f$ in $\Omega$. For $\varepsilon$ small enough, we can thus extend $\zeta$ in a tubular neighborhood of $\sigma_{1}(-\varepsilon,\varepsilon)$ foliated by the geodesic fibers $\exp_{x}(t\zeta|_{x})$, $x\in D$. In particular, this gives a smooth extension of $\zeta$ to $)\sigma_{2}([0,\tilde{\varepsilon}$)), for $\tilde{\varepsilon}$ small enough. Consider the map $F:[0,\tilde{\varepsilon})\times(-\varepsilon,\varepsilon)\to\mathbb{S}^{3}$ given by
$F(s,t)\doteq\exp_{\exp_{p}(s\sigma^{\prime}_{2}(0))}(t\zeta).$
[…]
By Claim 5 the extension of $\zeta$ provided in Claim 4 is in fact an extension to the entire $\mathbb{S}^{3}$. This concludes the proof of the lemma. ∎
Let us consider the quotient map $\pi:\mathbb{S}^{3}\to\mathbb{S}^{2}$ generated by $\zeta$, which by Lemma 3.3 is a smooth submersion. We claim that there exists a smooth map $g:\mathbb{S}^{2}\to\mathbb{S}^{2}$ such that
$f=g\circ\pi.$ (5.4)
The map $g$ is constructed as follows. Given $x\in\mathbb{S}^{2}$, consider the great circle $c_{x}\doteq\pi^{-1}(x).$ Since a fiber of $\pi$ is a connected component of a fiber of $f$, the circle $c_{x}$ is mapped by $f$ to a point $\tilde{x}=f(c_{x})$. We then define
$g(x)\doteq\tilde{x}$.
Since $\pi$ is a submersion, locally around $x$ it admits a smooth section $\sigma:V\subset\mathbb{S}^{2}\to\mathbb{S}^{3}$, and thus $g\equiv f\circ\sigma$ is smooth. Moreover, $g$ has rank $2$ at every point of the open set $\pi(U)$, in particular, $g$ is a global diffeomorphism if $f$ is a submersion. From Theorem 3.4 it follows that $\pi$ is a Hopf fibration, and it is immediate that $g$ must be conformal.
6. Proofs of the main theorems
6.1. Proof of Theorem A
Observe that the function $u^{2}-4v^{2}=(\lambda^{2}-\mu^{2})^{2}$ is real analytic and vanishes precisely at points where $\lambda=\mu$. Hence, if $\lambda=\mu$ in an open neighborhood, then $\lambda=\mu$ everywhere on $\mathbb{S}^{3}$. Assume now to the contrary that the continuous function
$w=(\lambda-\mu)^{2}\geq 0$
is not identically zero on $\mathbb{S}^{3}$. Notice that if $v^{2}=\lambda^{2}\mu^{2}=0$ on an open subset of $\mathbb{S}^{3}$, then it is zero everywhere by analyticity and from Lemma 4.3 we deduce that $f$ is constant. Therefore we may assume that the open set
$U\doteq\{x\in\mathbb{S}^{3}:\lambda(x)>\mu(x)>0\}$
is non-empty, and observe that $w$ is smooth on $U$. Assume that $w$ attains its maximum at a point $p_{0}$. If $p_{0}\in U$, then by Lemma 4.5 we have that
$b_{1k}^{4}(p_{0})=b_{2k}^{5}(p_{0}),\quad\text{for}\,\,\,k\in\{1,2,3\}.$ (6.1)
Moreover, the Laplacian of $w$ on $U$ is given by
$\frac{1}{2}\Delta w$ $=w(\lambda\mu+2)+|B|^{2}-2\sum_{k=1}^{3}(b_{1k}^{4}b_{2k}^{5}-b_{1k}^{5}b_{2k}^{4})-\frac{\mu}{\lambda}\sum_{k=1}^{3}(b_{3k}^{4})^{2}-\frac{\lambda}{\mu}\sum_{k=1}^{3}(b_{3k}^{5})^{2}.$ (6.2)
$=w(\lambda\mu+2)+(b_{11}^{4}-b_{12}^{5})^{2}+(b_{12}^{4}-b_{22}^{5})^{2}+(b_{11}^{5}+b_{12}^{4})^{2}+(b_{12}^{5}+b_{22}^{4})^{2}$
$\quad+\Big(2-\frac{\mu}{\lambda}\Big)\Big\{(b_{13}^{4})^{2}+(b_{23}^{4})^{2}\Big\}+\Big(2-\frac{\lambda}{\mu}\Big)\Big\{(b_{13}^{5})^{2}+(b_{23}^{5})^{2}\Big\}$
$\quad-2b_{13}^{4}b_{23}^{5}+2b_{13}^{5}b_{23}^{4}+\Big(1-\frac{\mu}{\lambda}\Big)(b_{33}^{4})^{2}+\Big(1-\frac{\lambda}{\mu}\Big)(b_{33}^{5})^{2}.$
By (6.1), harmonicity and Young’s inequality we obtain that at $p_{0}\in U$,
$\frac{1}{2}\Delta w\geq w(\lambda\mu+2)+\Big(2-\frac{\mu}{\lambda}-\frac{\lambda}{\mu}\Big)\Big(|B_{13}|^{2}+|B_{23}|^{2}+|B_{33}|^{2}\Big)\geq\Big(\lambda\mu+2-\frac{|B|^{2}}{\lambda\mu}\Big)w>0,$ (6.3)
which is a contradiction. Since we have assumed that $w$ is not identically zero, we have
$\lambda(p_{0})>\mu(p_{0})=0.$ (6.4)
(1) Let us examine the case
$|B|^{2}\leq\lambda\mu(\lambda\mu+a).$ (6.5)
In particular this implies that $|B|$ tends to zero as $\mu$ approaches zero. The set
$V\doteq\{\mu=0\}=\{v^{2}=0\}\subset\mathbb{S}^{3}$
is an analytic set of measure zero, and by our pinching it follows that
$|\nabla w|^{2}\leq C(\lambda-\mu)^{2}|B|^{2}\leq C(\lambda-\mu)^{2}\lambda\mu(\lambda\mu+a),$
whereby $\nabla w$ can be continuously extended by zero on $V$. It follows that $w$ is locally constant on $V$. By Remark 5.4, up to slightly moving $p_{0}$ we can assume that $p_{0}$ is a regular point of $V$. In particular, regardless of the dimension of $p_{0}$ as a regular point, there exists a small ball $D\subset U$ with $p_{0}\in\partial D$. Now from (6.3) and (6.5), we get that
$\begin{array}[]{lcl}\frac{1}{2}\Delta w&\geq&w(\lambda\mu+2)+2|B|^{2}-(\lambda^{2}+\mu^{2})(\lambda\mu+\alpha)\\[8.5359pt] &\longrightarrow&(2-\alpha)\lambda^{2}(p_{0})>0\qquad\text{as }\,p\to p_{0}.\end{array}$
For this reason, up to shrinking $D$ we can assume that
$\Delta w>0,\quad\text{on}\,\,\,D.$
If $\xi$ is the unit normal to the boundary of $D$ at $p_{0}$, then by Hopf’s Boundary Point Lemma, see for example [protter], we get that
$0=\langle\nabla w,\xi\rangle=\frac{\partial w}{\partial\xi}>0,$
which is a contradiction unless $w$ is a positive constant on $D$. However, in this latter case, from (6.3) we arrive again at contradiction. Hence $w=0$ and the result follows from Theorem 5.1.
(2) This pinching condition implies that $\lambda\mu$ is positive everywhere on $\mathbb{S}^{3}$, therefore $\mu$ cannot be zero which contradicts (6.4). Thus $w=0$ and the result follows from Theorem 5.1.
6.2. Proof of Theorem B
We proceed exactly as in the proof of Theorem A. Note that, under our assumptions, $B_{33}=0$ on $U$. Applying Young’s inequality to equation (6.2) we get
$\frac{1}{2}\Delta w\geq w(\lambda\mu+2)+\Big(2-\frac{\mu}{\lambda}-\frac{\lambda}{\mu}\Big)\Big(|B_{13}|^{2}+|B_{23}|^{2}\Big)\geq\Big(\lambda\mu+2-\frac{|B|^{2}}{2\lambda\mu}\Big)w\geq 0,$ (6.6)
this time on the entirety of $U$. Using Hopf’s Boundary Point Lemma, as in Theorem A, we get that $w$ must be constant on $\mathbb{S}^{3}$. Now by (6.6) we have that
$\left(\lambda\mu+2-\frac{|B|^{2}}{2\lambda\mu}\right)w\equiv 0,$
which, since we have assumed that $w$ is not identically zero, implies that
$|B|^{2}\equiv 2\lambda\mu(\lambda\mu+2).$
Then by Lemma 4.3 we get that
$\frac{1}{2}\Delta u=2(\lambda^{2}+\mu^{2})+4\lambda\mu=2(\lambda+\mu)^{2}\geq 0,$
and by the maximum principle $f$ must be constant, which is again a contradiction. Thus $w=0$ and the result follows from Theorem 5.1.
6.3. Proof of Theorem C
Denote as usual by $\lambda\geq\mu$ the two largest singular values. Assume that the open set $U=\{\lambda>\mu\}$ is non-empty. By Lemma 4.2, $\lambda^{2}$ and $\mu^{2}$ are smooth on $U$. Moreover, by assumption
$W(\lambda,\mu)=0,\qquad\mu W_{x}(\lambda,\mu)+\lambda W_{y}(\lambda,\mu)\neq 0\quad\text{on $U$}.$ (6.7)
Claim 1: The singular value $\mu$ is $C^{1}$ on $U$.
Proof of Claim 1. Differentiability may fail on the closed analytic subset
$\Sigma=\{\lambda>\mu=0\}\subset U,$
(the zero set of the analytic function $v^{2}$ in (2.4), see Lemma 4.3), on which in particular $\lambda$ is smooth; see [nomizu]. Differentiating (6.7) away from $\Sigma$ we get that
$W_{x}(\lambda,\mu)\,\nabla\lambda+W_{y}(\lambda,\mu)\nabla\mu=0.$
Since $W$ is $C^{1}$ and $\lambda$ and $\mu$ are continuous, the composition $W_{y}(\lambda,\mu)$ is continuous up to $\Sigma$, and the constraint in (6.7) forces it to be non-zero in the entire $U$. This implies the identity
$\nabla\mu=-\frac{W_{x}(\lambda,\mu)}{W_{y}(\lambda,\mu)}\,\nabla\lambda\qquad\text{on }\,U\backslash\Sigma.$
[…]
Claim 2: The continuous function $\varrho:\mathbb{S}^{3}\to\mathbb{R}$ given by
$\varrho=\lambda^{2}-\mu^{2}\geq 0$
is identically zero on $\mathbb{S}^{3}$.
Proof of Claim 2. Suppose to the contrary that this is not the case. Then the open set $U$ from Claim 1 is non-empty, and by Lemma 4.2 the function $\varrho$ is smooth and attains its maximum at a point $p_{0}\in U$. Differentiating $\varrho$ and using the relation (6.7) at $p_{0}$ we get that
$\begin{cases}W_{x}\,\alpha_{k}(\lambda)+W_{y}\,\alpha_{k}(\mu)=0,\\ \lambda\alpha_{k}(\lambda)-\mu\alpha_{k}(\mu)=0,\end{cases}$
whereby, by the constraint in (6.7), we obtain that at $p_{0}$
$\alpha_{k}(\lambda)=\alpha_{k}(\mu)=0.$ (6.8)
Equivalently, by Lemma 4.1, at $p_{0}$ we have that
$b_{11}^{4}=b_{12}^{4}=b_{13}^{4}=b_{12}^{5}=b_{22}^{5}=b_{23}^{5}=0,$
which in particular also implies that
$b_{11}^{5}=-b_{33}^{5}\quad\text{and}\quad b_{22}^{4}=-b_{33}^{4}.$
Substituting into the equation from Lemma 4.6 we immediately obtain that
$\Delta\varrho(p_{0})\geq 4\varrho(p_{0})>0,$
[…]
By Claim 2, $f$ is a weakly conformal harmonic map. Hence, by Theorem 5.1, the map $f$ is the composition of the Hopf fibration with a conformal map of $\mathbb{S}^{2}$.
6.4. Proof of Theorem D
First assume that the $2$-dilation of $f$ is a constant $c$. Then $f$ satisfies the assumptions in Theorem C with
W(x,y)=xyc.W(x,y)=xy-c.
Hence $f$ is weakly conformal, i.e. $\lambda\equiv\mu$ everywhere on $\mathbb{S}^{3}$. From $W(\lambda,\mu)=0$ we therefore deduce that $\lambda^{2}$ is constant. Escobales’ result [escobales] guarantees that $f$ is a Hopf fibration.
Next assume that the energy density $u$ is a positive constant $c$. From Lemma 4.3 we get
$|B|^{2}=2(v^{2}-u).$ (6.9)
Observe that $\mu$ cannot vanish at any point $p$, otherwise $v$ would vanish at $p$ and thus from (6.9) we get $u=|B|^{2}=0$ at $p$, contradiction. Consequently, $f$ is a submersion. Setting
$W(x,y)=x^{2}+y^{2}-c,$
even though (1.2) is not satisfied on $\{x>y=0\}$, nevertheless
$\mu W_{x}(\lambda,\mu)+\lambda W_{y}(\lambda,\mu)\neq 0\qquad\text{holds everywhere on }\,\mathbb{S}^{3}.$
This suffices to follow the proof of Theorem C and conclude that $f$ is weakly conformal. From $\lambda\equiv\mu$ and $u=\lambda^{2}+\mu^{2}$ constant we deduce that $\lambda$ and $\mu$ are constant. The conclusion follows by Escobales’ result.